# Towards an algebraic approach to the reconfiguration CSP

### v1

@@ 2 Preliminaries
In this section, we first define the Reconfiguration CSP (RCSP). Then, in Section 2.3, we introduce partial operations and explain the reduction relationships between RCSPs. Furthermore, in Section 2.4, we summarize known results in the context of Boolean RCSPs.
@@ 2.3 Partial polymorphism and logical expression (reduction)
A $k$-ary partial operation $f$ on $D$ is a mapping $f:D^{\prime}\rightharpoonup D$ for some $D^{\prime}\subseteq D^{k}$. Here, $D^{\prime}$ is called the domain of $f$ and is denoted by ${\rm dom}(f)$. If $\boldsymbol{x}\in D^{\prime}$, then we say that $f(\boldsymbol{x})$ is defined and otherwise (that is, if $\boldsymbol{x}\notin D^{\prime}$) undefined. If ${\rm dom}(f)=D^{k}$, then $f$ is a total operation.
@@
A result similar to proposition 2.14 is known.
@@
This follows from proposition 2.25, since $s(I)$ can be represented as a $\varGamma$-formula. ∎
@@
Given an instance $(I,\boldsymbol{s},\boldsymbol{t})$ of ${\rm RCSP}(\varGamma_{2})$ with $n$ variables, we transform it into an equivalent instance $(I^{\prime},\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$ of ${\rm RCSP}(\varGamma_{1})$. From proposition 2.25, every constraint $R(x_{1},...,x_{r})$ in $I$ can be replaced with constraints
@@ 2.4 Known results for the Boolean RCSP
Recall the example in Example 2.5. There, $R^{\prime}_{01}(x_{2}):=R_{01}(0,x_{2})$ is obtained from $R_{01}$ by substitution of constants. Moreover, $R^{\prime}_{00}(x_{3}):=R_{00}(x_{3},x_{3})$ is obtained from $R_{00}$ by identification of variables.
@@ 3.1 Ordered partial Maltsev operation and characterization of $\varGamma_{{\rm SOF}}$
By definition, an ordered partial Maltsev operation is obtained by restricting the domain of the partial Maltsev operation. Note that by definition ordered partial Maltsev operations are idempotent.
@@
When $D=\{0,1\}$ and $0<1$, the $(D,\leq)$-Maltsev operation $M_{D,\leq}$ is a ternary operation $M_{D,\leq}:\{0,1\}^{3}\rightharpoonup\{0,1\}$ such that $M(0,0,0)=M(0,1,1)=M(1,1,0)=0$, $M(1,1,1)=1$, and undefined for the other inputs.
@@
When $D=\{0,1\}$ and $1<0$, the $(D,\leq)$-Maltsev operation $M_{D,\leq}$ is a ternary operation $M_{D,\leq}:\{0,1\}^{3}\rightharpoonup\{0,1\}$ such that $M(0,0,0)=0$, $M(1,0,0)=M(0,0,1)=M(1,1,1)=1$, and undefined for the other inputs.
We will show that the partial operations in examples 3.2 and 3.3 characterize the sets of safely OR-free and safely NAND-free relations, respectively.
@@
Let $D=\{0,1\}$ and $0<1$. Let $r$ be a positive integer and $R$ is an $r$-ary relation $R\subseteq\{0,1\}^{r}$. Then, $R$ is safely OR-free if and only if it is invariant under the $(D,\leq)$-Maltsev operation.
@@ 3.2 Algorithm for RCSP invariant under an ordered partial Maltsev operation
Now we extend the algorithmic result for ${\rm RCSP}_{\rm C}(\varGamma_{{\rm SOF}})$ by Gopalan et al. [GKMP09] to the case where $\varGamma$ is invariant under an ordered partial Maltsev operation on an arbitrary finite domain. The algorithm is similar to the one for ${\rm RCSP}_{\rm C}(\varGamma_{{\rm SOF}})$ in [GKMP09]. First, we show that there is a unique locally minimal solution for each connected component of the solution graph of any instance of ${\rm RCSP}(\varGamma)$. Here, a solution is locally minimal in the solution graph if it has no smaller neighboring element in the solution graph. This shows that by greedily descending from any solution to its neighbors, we can efficiently find the unique locally minimal solution of the connected component to which the solution belongs. Thus, the reconfiguration problem can be solved by finding the unique minimal solution of the connected component to which each solution belongs for two given solutions and checking whether they coincide.
@@
Let $\boldsymbol{u}=M_{D,\leq}(\boldsymbol{u}^{i-1},\boldsymbol{u}^{i},\boldsymbol{u}^{i+1})$. Then it holds that $u_{1}=p,u_{2}=s$, and $u_{j}=u^{i}_{j}$ for $j>2$. Since $s(I)$ is invariant under $M_{D,\leq}$ by corollary 2.26, $\boldsymbol{u}$ is also a solution of $I$ and the path $\boldsymbol{u}^{0}\rightarrow\boldsymbol{u}^{1}\rightarrow\dots\rightarrow\boldsymbol{u}^{i-1}\rightarrow\boldsymbol{u}\rightarrow\boldsymbol{u}^{i+1}\rightarrow\dots\rightarrow\boldsymbol{u}^{\ell}$ is also an $(\boldsymbol{a},\boldsymbol{b})$-path in $C$. However, this contradicts the way we chose the original path $P$, since $\boldsymbol{u}^{i-1}\geq\boldsymbol{u}$. It follows that $C$ contains only one locally minimal element.
@@
Let $(I,\boldsymbol{s},\boldsymbol{t})$ be an instance of ${\rm RCSP}_{\rm C}(\varGamma)$. From lemma 3.7 we can compute a unique minimal element $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) of the connected component that contains $\boldsymbol{s}$ (resp, $\boldsymbol{t}$) by greedily continuing to decreasing the solution as far as possible. Determining if some value of a variable can be decreased is done in ${\rm O}(nm|D|)$ time. Since the value of each variable can be decreased at most |D|1|D|-1 times, we can reach to $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) by decreasing a value of some variable at most n(|D|1)n(|D|-1) times. Therefore, $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) can be obtained in ${\rm O}(n^{2}m|D|^{2})$ time. We output yes if $\boldsymbol{s}_{\min}=\boldsymbol{t}_{\min}$ and no otherwise. ∎
@@ 3.3 Relations invariant under some ordered partial Maltsev operation
General
CSPs
Min-closed [Thm 3.13]
Strongly rectangular [DR10]
Lin. eqs. over GF(qq[JCG95]
Boolean
CSPs
Safely OR-free [Thm 3.5]
Safely NAND-free [Thm 3.6]
Exact SAT [LW21]
Subset sum [LW21]
Lin. eqs. over ${\rm GF}(2)$ [Sch78]
Graphs Rectangular [Thm 3.17] Circular clique $C_{6,3}$ [Prop 3.24]
Digraphs
Trans. tourn. [Thm 3.26]
$r(\overrightarrow{K_{n}})$ [Thm 3.26]
Rectangular [Thm 3.17]
Totally rectangular [CEJN15]
Graphs in Fig. 2 [Prop 3.23]
@@ 3.3.1 Relations invariant under some ordered partial Maltsev operation
Note that for any Maltsev operation $M$ and the partial Maltsev operation $M_{p}$ on $D$, and any total order $\leq$ on $D$, we have $M_{D,\leq}\leq M_{p}\leq M$ (for the order relations between partial operations, please refer to Definition 2.20). Thus, if a constraint language $\varGamma$ is invariant under a Maltsev operation, it is also invariant under a partial Maltsev operation, and further under some ordered partial Maltsev operation by Lemma 2.23.
@@ 3.3.2 Digraphs invariant under some ordered partial Maltsev operation
By theorem 3.16 it suffices to show that the 4-cycle is totally rectangular. It can be verified that for each pair of diagonally opposite vertices (say, $u$ and $v$) of the 4-cycle, we have $u^{+k}=v^{+k}$. Moreover, for distinct vertices $u$ and $v$ that are not diagonally opposite, we have $x^{+k}\cap y^{+k}=\emptyset$. Therefore, the 4-cycle is $k$-rectangular by lemma 3.19. Since $k$ is an arbitrary positive integer, the 4-cycle is totally rectangular. ∎
@@
The vertex set of $C_{6.3}$ is $\{0,1,\dots,5\}$ and the edge set of it is $\{03,14,25\}$. It can be verified that for any distinct $u,v\in V(C_{6,3})$ and any positive integer $k$, we have $u^{+k}\cap v^{+k}=\emptyset$. Hence, by lemma 3.19, $C_{6.3}$ is totally rectangular, and thus invariant under a Maltsev operation by theorem 3.16. ∎
@@ 4 Safely componentwise bijunctive relations
Now, since $C_{1}$ and $C_{2}$ are invariant and the majority operation on $\{0,1\}$ by lemma 2.32, $C_{1}\times C_{2}$ is also invariant under the majority operation. Moreover, every connected component of a relation invariant under the majority operation is also invariant under the majority operation by lemma 2.17. Furthermore, identification of variables preserves invariance under an operation. Hence, $C$ is also invariant under the majority operation. This means that $C$ is bijunctive by lemma 2.32. Therefore, $R$ is safely componentwise bijunctive.
Now, the latter statement follows from Proposition 2.25. ∎
@@
First, we show that $R$ is componentwise bijunctive. Let $C$ be a connected component of $R$. From lemma 2.32 in Subsection 2.4, it suffices to show that for any $\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3}\in C$ we have ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})\in C$. Take any tuples $\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3}\in C$. By assumption, these tuples are comparable under $\leq_{o}$. Assume without loss of generality that $\boldsymbol{t}^{1}\leq_{o}\boldsymbol{t}^{2}\leq_{o}\boldsymbol{t}^{3}$. We claim that ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})=\boldsymbol{t}^{2}$. Indeed, for any index $j\in[r]$, $t^{2}_{j}$ is the median among the values $t^{1}_{j}$, $t^{2}_{j}$, and $t^{3}_{j}$. Thus, ${\rm maj}(t^{1}_{j},t^{2}_{j},t^{3}_{j})=t^{2}_{j}$ for every $j\in[r]$ and ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})=\boldsymbol{t}^{2}$ holds. Therefore, ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})\in C$.
@@
First, we show that $M^{(r)}$ is not safely componentwise bijunctive. It is easy to see that $G(M^{(r)})$ is connected as ${\rm dist}(\boldsymbol{u}^{r,i},\boldsymbol{u}^{r,i+1})=1$ for every $i\in[r+1]$ by construction. Moreover, we claim that $\boldsymbol{u}={\rm maj}(\boldsymbol{u}^{r,1},\boldsymbol{u}^{r,3},\boldsymbol{u}^{r,r+2})$ is not in $M^{(r)}$. Indeed, $\boldsymbol{u}^{r,1}$ and $\boldsymbol{u}^{r,3}$ differ only in the first and second coordinates. Therefore, $u_{j}=u^{r,1}_{j}$ for all $j\geq 3$. Moreover, we have $u^{r,1}_{1}=u^{r,r+2}_{1}=0$ and $u^{r,3}_{2}=u^{r,r+2}_{2}=1$, implying that $(u_{1},u_{2})=(0,0)$. Since $(u^{r,i}_{1},u^{r,i}_{2})\neq(0,0)$ for all $i\in[r+1]$ and $u_{3}=0\neq 1=u^{r,r+2}_{3}$, we conclude that $\boldsymbol{u}$ is not in $M^{(r)}$. Therefore, by lemma 2.32 in Subsection 2.4, $M^{(r)}$ is not componentwise bijunctive and thus not safely componentwise bijunctive.
@@
Next, we show that for any $R\subsetneq M^{(r)}$ it holds that $R$ is safely componentwise bijunctive. Let $R\subsetneq M^{(r)}$. It is easy to see that each connected component of $R$ admits a total order. Moreover, let us assume that $R$ has two or more connected components. Assume that $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$, where $1\leq i<k\leq r+2$, are in different connected components, say $C_{1}$ and $C_{2}$. We will show that these tuples are never in the same connected component even after identification of variables is applied to $R$. Indeed, since $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$ are disconnected, there exists $p\in\{i+1,\dots,k-1\}$ such that $\boldsymbol{u}^{r,p}$ is not in $R$. It follows that $k-i\geq 2$. Now, we show that there exist coordinates $j$ and $j^{\prime}$ such that $u^{r,i}_{j}=0$ and $u^{r,k}_{j}=1$, and $u^{r,i}_{j^{\prime}}=1$ and $u^{r,k}_{j^{\prime}}=0$. To show this, assume first that $k\neq r+2$. If $p$ is even, then we see that $u^{r,i}_{p-1}=0$ and $u^{r,k}_{p-1}=1$, and $u^{r,i}_{p}=1$ and $u^{r,k}_{p}=0$. If $p$ is odd, then we see that $u^{r,i}_{p-1}=1$ and $u^{r,k}_{p-1}=0$, and $u^{r,i}_{p}=0$ and $u^{r,k}_{p}=1$, as desired. The case of $k=r+2$ and $i\neq 1$ can be proven similarly. Finally, if $i=1$ and $k=r+2$, then $u^{r,1}_{2}=1$ and $u^{r,r+2}_{2}=0$, and $u^{r,1}_{3}=0$ and $u^{r,r+2}_{3}=1$, as desired. Now, we show that $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$ are never in the same connected component even after identification of variables is applied to $R$. This is because, even if identification of variables is applied to $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$, $u^{r,i}_{j}=0$ and $u^{r,k}_{j}=1$, and $u^{r,i}_{j^{\prime}}=1$ and $u^{r,k}_{j^{\prime}}=0$ are preserved. Hence, the distances between the tuples in $C_{1}$ and $C_{2}$ are at least two and they remain disconnected. Hence, after the application of variable identification to $R$, each connected component still admits a total order as shown in the proof of lemma 4.3. Therefore, it is also componentwise bijunctive. ∎
@@
This theorem follows from Lemma 4.4 and the argument described below the theorem. ∎

### v2

@@ 2 Preliminaries
In this section, we first define the Reconfiguration CSP (RCSP). Then, in Subsection 2.3, we introduce partial operations and explain the reduction relationships between RCSPs. Furthermore, in Subsection 2.4, we summarize known results in the context of Boolean RCSPs.
@@ 2.3 Partial polymorphism and logical expression (reduction)
A $k$-ary partial operation $f:D^{k}\rightharpoonup D$ on $D$ is a mapping $f:D^{\prime}\to D$ for some $D^{\prime}\subseteq D^{k}$. Here, $D^{\prime}$ is called the domain of $f$ and is denoted by ${\rm dom}(f)$. If $\boldsymbol{x}\in D^{\prime}$, then we say that $f(\boldsymbol{x})$ is defined and otherwise (that is, if $\boldsymbol{x}\notin D^{\prime}$) undefined. If ${\rm dom}(f)=D^{k}$, then $f$ is a total operation.
@@
A result similar to Subsection 2.2 is known.
@@
This follows from Subsection 2.3, since $s(I)$ can be represented as a $\varGamma$-formula. ∎
@@
Given an instance $(I,\boldsymbol{s},\boldsymbol{t})$ of ${\rm RCSP}(\varGamma_{2})$ with $n$ variables, we transform it into an equivalent instance $(I^{\prime},\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$ of ${\rm RCSP}(\varGamma_{1})$. From Subsection 2.3, every constraint $R(x_{1},...,x_{r})$ in $I$ can be replaced with constraints
@@ 2.4 Known results for the Boolean RCSP
Recall the example in Subsection 2.1. There, $R^{\prime}_{01}(x_{2}):=R_{01}(0,x_{2})$ is obtained from $R_{01}$ by substitution of constants. Moreover, $R^{\prime}_{00}(x_{3}):=R_{00}(x_{3},x_{3})$ is obtained from $R_{00}$ by identification of variables.
@@ 3.1 Ordered partial Maltsev operation and characterization of $\varGamma_{{\rm SOF}}$
By definition, an ordered partial Maltsev operation is a subfunction of the partial Maltsev operation. Note that by definition ordered partial Maltsev operations are idempotent.
@@
When $D=\{0,1\}$ and $0<1$, the $(D,\leq)$-Maltsev operation $M_{D,\leq}$ is a ternary partial operation $M_{D,\leq}:\{0,1\}^{3}\rightharpoonup\{0,1\}$ such that $M(0,0,0)=M(0,1,1)=M(1,1,0)=0$, $M(1,1,1)=1$, and undefined for the other inputs.
@@
When $D=\{0,1\}$ and $1<0$, the $(D,\leq)$-Maltsev operation $M_{D,\leq}$ is a ternary partial operation $M_{D,\leq}:\{0,1\}^{3}\rightharpoonup\{0,1\}$ such that $M(0,0,0)=0$, $M(1,0,0)=M(0,0,1)=M(1,1,1)=1$, and undefined for the other inputs.
We will show that the partial operations in Subsections 3.1 and 3.1 characterize the sets of safely OR-free and safely NAND-free relations, respectively.
@@
Let $D=\{0,1\}$ and $0<1$. Let $r$ be a positive integer and $R$ be an $r$-ary relation $R\subseteq\{0,1\}^{r}$. Then, $R$ is safely OR-free if and only if it is invariant under the $(D,\leq)$-Maltsev operation.
@@ 3.2 Algorithm for RCSP invariant under an ordered partial Maltsev operation
Now we extend the algorithmic result for ${\rm RCSP}_{\rm C}(\varGamma_{{\rm SOF}})$ by Gopalan et al. [GKMP09] to the case where $\varGamma$ is invariant under an ordered partial Maltsev operation on an arbitrary finite domain. The algorithm is similar to the one for ${\rm RCSP}_{\rm C}(\varGamma_{{\rm SOF}})$ in [GKMP09]. First, we show that there is a unique locally minimal solution for each connected component of the solution graph of any instance of ${\rm RCSP}(\varGamma)$. Here, a solution is locally minimal in the solution graph if it has no smaller neighboring element in the solution graph. This shows that by greedily descending from any solution to its neighbors, we can efficiently find the unique locally minimal solution of the connected component to which the solution belongs. Thus, the reconfiguration problem can be solved by finding the unique locally minimal solution of the connected component to which each solution belongs for two given solutions and checking whether they coincide.
@@
Let $\boldsymbol{u}=M_{D,\leq}(\boldsymbol{u}^{i-1},\boldsymbol{u}^{i},\boldsymbol{u}^{i+1})$. Then it holds that $u_{1}=p,u_{2}=s$, and $u_{j}=u^{i}_{j}$ for $j>2$. Since $s(I)$ is invariant under $M_{D,\leq}$ by Subsection 2.3, $\boldsymbol{u}$ is also a solution of $I$ and the path $\boldsymbol{u}^{0}\rightarrow\boldsymbol{u}^{1}\rightarrow\dots\rightarrow\boldsymbol{u}^{i-1}\rightarrow\boldsymbol{u}\rightarrow\boldsymbol{u}^{i+1}\rightarrow\dots\rightarrow\boldsymbol{u}^{\ell}$ is also an $(\boldsymbol{a},\boldsymbol{b})$-path in $C$. However, this contradicts the way we chose the original path $P$, since $\boldsymbol{u}^{i-1}\geq\boldsymbol{u}$. It follows that $C$ contains only one locally minimal element.
@@
Let $(I,\boldsymbol{s},\boldsymbol{t})$ be an instance of ${\rm RCSP}_{\rm C}(\varGamma)$. From Subsection 3.2 we can compute a unique minimal element $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) of the connected component that contains $\boldsymbol{s}$ (resp, $\boldsymbol{t}$) by greedily continuing to decreasing the solution as far as possible. Determining if some value of a variable can be decreased is done in ${\rm O}(nm|D|)$ time. Since the value of each variable can be decreased at most |D|1|D|-1 times, we can reach to $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) by decreasing a value of some variable at most n(|D|1)n(|D|-1) times. Therefore, $\boldsymbol{s}_{\min}$ (resp., $\boldsymbol{t}_{\min}$) can be obtained in ${\rm O}(n^{2}m|D|^{2})$ time. We output yes if $\boldsymbol{s}_{\min}=\boldsymbol{t}_{\min}$ and no otherwise. ∎
@@ 3.3 Relations invariant under some ordered partial Maltsev operation
General CSPs Min-closed [Thm 3.3.1] Strongly rectangular [DR10] Lin. eqs. over GF(qq[JCG95]
Boolean CSPs Safely OR-free [Thm 3.5] Safely NAND-free [Thm 3.6] Exact SAT [LW21] Subset sum [LW21] Lin. eqs. over ${\rm GF}(2)$ [Sch78]
Graphs Rectangular [Thm 3.17] Circular clique $C_{6,3}$ [Prop 3.3.2]
Digraphs Trans. tourn. [Thm 3.26] $r(\overrightarrow{K_{n}})$ [Thm 3.26] Rectangular [Thm 3.17] Totally rectangular [CEJN15] Graphs in Fig. 2 [Prop 3.3.2]
@@ 3.3.1 Relations invariant under some ordered partial Maltsev operation
Note that for any Maltsev operation $M$ and the partial Maltsev operation $M_{p}$ on $D$, and any total order $\leq$ on $D$, we have $M_{D,\leq}\leq M_{p}\leq M$ (for the order relations between partial operations, please refer to Subsection 2.3). Thus, if a constraint language $\varGamma$ is invariant under a Maltsev operation, it is also invariant under a partial Maltsev operation, and further under some ordered partial Maltsev operation by Subsection 2.3.
@@ 3.3.2 Digraphs invariant under some ordered partial Maltsev operation
By Theorem 3.16 it suffices to show that the 4-cycle is totally rectangular. It can be verified that for each pair of diagonally opposite vertices (say, $u$ and $v$) of the 4-cycle, we have $u^{+k}=v^{+k}$. Moreover, for distinct vertices $u$ and $v$ that are not diagonally opposite, we have $x^{+k}\cap y^{+k}=\emptyset$. Therefore, the 4-cycle is $k$-rectangular by Subsection 3.3.2. Since $k$ is an arbitrary positive integer, the 4-cycle is totally rectangular. ∎
@@
The vertex set of $C_{6.3}$ is $\{0,1,\dots,5\}$ and the edge set of it is $\{03,14,25\}$. It can be verified that for any distinct $u,v\in V(C_{6,3})$ and any positive integer $k$, we have $u^{+k}\cap v^{+k}=\emptyset$. Hence, by Subsection 3.3.2, $C_{6.3}$ is totally rectangular, and thus invariant under a Maltsev operation by Theorem 3.16. ∎
@@ 4 Safely componentwise bijunctive relations
Now, since $C_{1}$ and $C_{2}$ are invariant and the majority operation on $\{0,1\}$ by Subsection 2.4, $C_{1}\times C_{2}$ is also invariant under the majority operation. Moreover, every connected component of a relation invariant under the majority operation is also invariant under the majority operation by Subsection 2.2. Furthermore, identification of variables preserves invariance under an operation. Hence, $C$ is also invariant under the majority operation. This means that $C$ is bijunctive by Subsection 2.4. Therefore, $R$ is safely componentwise bijunctive.
Now, the latter statement follows from Subsection 2.3. ∎
@@
First, we show that $R$ is componentwise bijunctive. Let $C$ be a connected component of $R$. From Subsection 2.4 in Subsection 2.4, it suffices to show that for any $\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3}\in C$ we have ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})\in C$. Take any tuples $\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3}\in C$. By assumption, these tuples are comparable under $\leq_{o}$. Assume without loss of generality that $\boldsymbol{t}^{1}\leq_{o}\boldsymbol{t}^{2}\leq_{o}\boldsymbol{t}^{3}$. We claim that ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})=\boldsymbol{t}^{2}$. Indeed, for any index $j\in[r]$, $t^{2}_{j}$ is the median among the values $t^{1}_{j}$, $t^{2}_{j}$, and $t^{3}_{j}$. Thus, ${\rm maj}(t^{1}_{j},t^{2}_{j},t^{3}_{j})=t^{2}_{j}$ for every $j\in[r]$ and ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})=\boldsymbol{t}^{2}$ holds. Therefore, ${\rm maj}(\boldsymbol{t}^{1},\boldsymbol{t}^{2},\boldsymbol{t}^{3})\in C$.
@@
First, we show that $M^{(r)}$ is not safely componentwise bijunctive. It is easy to see that $G(M^{(r)})$ is connected as ${\rm dist}(\boldsymbol{u}^{r,i},\boldsymbol{u}^{r,i+1})=1$ for every $i\in[r+1]$ by construction. Moreover, we claim that $\boldsymbol{u}={\rm maj}(\boldsymbol{u}^{r,1},\boldsymbol{u}^{r,3},\boldsymbol{u}^{r,r+2})$ is not in $M^{(r)}$. Indeed, $\boldsymbol{u}^{r,1}$ and $\boldsymbol{u}^{r,3}$ differ only in the first and second coordinates. Therefore, $u_{j}=u^{r,1}_{j}$ for all $j\geq 3$. Moreover, we have $u^{r,1}_{1}=u^{r,r+2}_{1}=0$ and $u^{r,3}_{2}=u^{r,r+2}_{2}=1$, implying that $(u_{1},u_{2})=(0,0)$. Since $(u^{r,i}_{1},u^{r,i}_{2})\neq(0,0)$ for all $i\in[r+1]$ and $u_{3}=0\neq 1=u^{r,r+2}_{3}$, we conclude that $\boldsymbol{u}$ is not in $M^{(r)}$. Therefore, by Subsection 2.4 in Subsection 2.4, $M^{(r)}$ is not componentwise bijunctive and thus not safely componentwise bijunctive.
@@
Next, we show that for any $R\subsetneq M^{(r)}$ it holds that $R$ is safely componentwise bijunctive. Let $R\subsetneq M^{(r)}$. It is easy to see that each connected component of $R$ admits a total order. Moreover, let us assume that $R$ has two or more connected components. Assume that $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$, where $1\leq i<k\leq r+2$, are in different connected components, say $C_{1}$ and $C_{2}$. We will show that these tuples are never in the same connected component even after identification of variables is applied to $R$. Indeed, since $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$ are disconnected, there exists $p\in\{i+1,\dots,k-1\}$ such that $\boldsymbol{u}^{r,p}$ is not in $R$. It follows that $k-i\geq 2$. Now, we show that there exist coordinates $j$ and $j^{\prime}$ such that $u^{r,i}_{j}=0$ and $u^{r,k}_{j}=1$, and $u^{r,i}_{j^{\prime}}=1$ and $u^{r,k}_{j^{\prime}}=0$. To show this, assume first that $k\neq r+2$. If $p$ is even, then we see that $u^{r,i}_{p-1}=0$ and $u^{r,k}_{p-1}=1$, and $u^{r,i}_{p}=1$ and $u^{r,k}_{p}=0$. If $p$ is odd, then we see that $u^{r,i}_{p-1}=1$ and $u^{r,k}_{p-1}=0$, and $u^{r,i}_{p}=0$ and $u^{r,k}_{p}=1$, as desired. The case of $k=r+2$ and $i\neq 1$ can be proven similarly. Finally, if $i=1$ and $k=r+2$, then $u^{r,1}_{2}=1$ and $u^{r,r+2}_{2}=0$, and $u^{r,1}_{3}=0$ and $u^{r,r+2}_{3}=1$, as desired. Now, we show that $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$ are never in the same connected component even after identification of variables is applied to $R$. This is because, even if identification of variables is applied to $\boldsymbol{u}^{r,i}$ and $\boldsymbol{u}^{r,k}$, $u^{r,i}_{j}=0$ and $u^{r,k}_{j}=1$, and $u^{r,i}_{j^{\prime}}=1$ and $u^{r,k}_{j^{\prime}}=0$ are preserved. Hence, the distances between the tuples in $C_{1}$ and $C_{2}$ are at least two and they remain disconnected. Hence, after the application of variable identification to $R$, each connected component still admits a total order as shown in the proof of Section 4. Therefore, it is also componentwise bijunctive. ∎
@@
This theorem follows from Section 4 and the argument described below the theorem. ∎